OpenStudy (anonymous):

Can you tell me how to write the equation for the horizontal line that contains point G(-8, 8)?

OpenStudy (nurali):

A horizontal line has no slope. So when calculating your equation of a line, you want to use point-slope form: (y - y1) = m(x - x1), where (x1,y1) is your coordinates. So... (y - 8) = 0 (x - (-8)) y - 8 = 0 y = 8 Now if you want to think about it, this line extends horizontal, so it runs parallel to the x-axis. It must go through that point, so it has to be y = 8 and only y = 8 for it to remain perfectly horizontal while going through that point.

OpenStudy (anonymous):

U mean like this? m = 0 x1 = -8 y1 = 8 y - 8 = 0(x - (-8)) y - 8 = 0 y = 8 Is that correct?

OpenStudy (nurali):

Yes

OpenStudy (anonymous):

Oh thank u sooooooo much! I think I got it :)

OpenStudy (nurali):

My Pleasure.

OpenStudy (anonymous):

So how will I give u a reward? This is ,y first time in open study....

OpenStudy (anonymous):

How do I get new question?